OpenStudy (sonyalee77):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line y=7/5x -3/5? A. 5x + 7y = 1 B. 7x + 5y = –13 C. 7x + 5y = –1 D. 3x + 5y = 7

OpenStudy (sonyalee77):

@Euler271

OpenStudy (anonymous):

the line is in the form y = mx + b, where m is the slope and b in the y-intercept. the slope in this case is 7/5. the slope of perpendicular lines are the NEGATIVE RECIPROCAL of eachother. so the slope of the line perpendicular to the line of slope m is (-1/m). so the slope of the line perpendicular to our line here is (-5/7). the y-intercept of the original line will not be used since it is irrelevant information in our case. to find the equation of the line given the slope and given a point (Xo, Yo), you take: \[(y - y_0) = m(x - x_0)\]\[y - 3 = \left(- \frac{ 5 }{ 7 } \right)(x - (-4))\]\[y = 3 - \frac{ 5 }{ 7 }x -\frac{ 20 }{ 7 }\]\[y = - \frac{ 5 }{ 7 } + \frac{ 1 }{ 7 }\]

OpenStudy (sonyalee77):

is it A.

OpenStudy (sonyalee77):

I think that it's A am I right?

OpenStudy (anonymous):

yes (my bad for delay lol)